OpenStudy (iluvsoccer):

What are the x and y-intercepts of the line described by this equation? 3x−9y=10.8 Enter your answers in the boxes in decimal form. x-intercept = y-intercept =

OpenStudy (iluvsoccer):

@SerenityJustice74

OpenStudy (serenityjustice74):

Just make them each zero 3(0) - 9y = 10.8 -y = 10.8/9 y = -1.2 3x - 9(0) = 10.8 3x = 10.8 x = 10.8/3 x= 3.6

OpenStudy (iluvsoccer):

Wow. That is easy. :o Thanks! :)

OpenStudy (serenityjustice74):

Mhm solving for intercepts is easy